OpenStudy (anonymous):

If f(x) = -2x + 1, find f(-6).

OpenStudy (anonymous):

can any one help ?

OpenStudy (anonymous):

f(6) = -2(-6)+1

OpenStudy (anonymous):

@AlexTra these are the choses they gave me If f(x) = -2x + 1, find f(-6). 12 13 -11 -7

OpenStudy (anonymous):

The answer is there do the math -2*(-6)+1

OpenStudy (anonymous):

so the answer would be -7 right @AlexTra

OpenStudy (anonymous):

13

OpenStudy (anonymous):

(-2)*(-6) =12+1 = 13

OpenStudy (anonymous):

how did you get 13?

OpenStudy (anonymous):

@AlexTra how did you get 13 ?